I have 19 coins
The number of dimes is three times the number of quarters
The number of nickels is one less than twice the number of quarters
The number of pennies is two more than dimes
How much money do I have

Answer by ikleyn(52803)

.

quarters + dimes + nickels + pennies = 19


x        + 3x    + (2x-1)  + (3x+2)  = 19


9x                                   = 19 + 1 - 2 = 18


x                                                 = 18/9 = 2.


ANSWER.  2 quarters,  3*2 = 6 dimes, 2*2-1 = 3 nickels and 3*2+2 = 8 pennies.


CHECK.  Total coins = 2 + 6 + 3 + 8 = 19.   ! Correct !

Solved.